OpenStudy (anonymous):

Rearrange this equation to isolate c a=b(1/c - 1/d)

OpenStudy (anonymous):

you have to first divide by b, add 1/d, then multiply by c. Then divide by everything that is not c.

OpenStudy (anonymous):

Did you get an answer yet(Im trying to find out where you are) or are u clueless?

OpenStudy (anonymous):

thanks..thats the part I was not getting.. the multiply by c

OpenStudy (anonymous):

a/b+1/d = 1/c c*[a/b+1/d] = 1 c = 1 / [a/b +1/d]

OpenStudy (anonymous):

thank you!